• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name E.Chudakov

    Log entry time 16:25:45 on April 02, 2008

    Entry number 229550

    This entry is a followup to: 229014

    keyword=Moller polarimeter status

    The collimator remote controller has been fixed. Two wires have been disconnected from the controller, Jack Segal plugged them back.

    The collimator is left at about 30mm opening (7000 steps from the closed position).

    We used the primitive telnet interface to the controller. The GUI program seems to be confused. At the "negative end switch", corresponding to the collimator closed, the program decides to flip the direction of motion and instead of opening, starts closing.



    A copy of this log entry has been emailed to: gorbenko@jlab.org, romanip@jlab.org